# Renormalization in Lorenz maps - completely invariant sets and periodic orbits

### v3

@@ 1. Introduction
The main object of our study are expanding Lorenz maps, that is maps $f\colon[0,1]\to[0,1]$ satisfying the following three conditions:
@@
with special emphasis on piecewise linear case. It turned out that such maps appear in a natural way as Poincarè maps in geometric models of well known Lorenz attractor. This description was independently appeared in works of Guckenheimer [21], Williams [37] and Afraimovich, Bykov and Shil‘nikov [4] as a tool for better understanding of chaotic behavior present in Lorenz model (cf. more recent [27]). In 1985, Hofbauer performed in [22] an analysis of possible periods of periodic points in Lorenz maps by introducing special type of infinite graphs, induced by evolution of pieces of monotonicity under iteration of Lorenz map. While a characterization similar to Scharkovsky was provided later in [3] by different methods, the tools developed by Hofbauer are of independent interest, since they provide a deep insight into dynamics of Lorenz maps. An extensive study on utility of Hofbauer’s approach, kneading theory of Milnor and Thurston and many other interesting results on dynamics of Lorenz maps can be found in PhD thesis of M. St. Pierre, see [32] (cf. [26]). Even the simplest case of maps $x\mapsto\beta x$ generates many difficulties and leads to deep results. It is usually performed at the level of symbolic dynamics, so-called $\beta$-shifts. In [31], Schmeling showed that many of these maps do not have the specification property, and later Thompson [36] provided nice representaions of these systems together with their further descriptions (see [12] for recent advances that originated from this, together with a nice survey on studies on structure of entropy and topological pressure). Mixing in $x\mapsto\beta x+\alpha$ also reveals complicated structure of parameters [33] (cf. [14] for measure-theoretic mixing in a similar family of maps).
@@
    (1.1) $R_{E}f(x)=\begin{cases}f^{l}(x)&,x\in[f^{r}(c_{+}),c)\\ f^{r}(x)&,x\in(c,f^{l}(c_{-})]\\ \end{cases}$
@@
Theorem 1.1 is aiming at providing a beautiful connection between renormalizations and completely invariant sets. We are going to show that unfortunately, this result is not true as stated, however we are able to provide some possible alternatives (c.f. Lemma 2.9 or Example 3.9). This way we show that some connections exists, however their complete description in full generality of Lorenz maps may be a challenging problem. At this point it is worth to mention closely related recent studies of R. Alcazar Barrera [1] (see also [2]), initiated in his PhD thesis, on invariant sets arising from omitting gaps is subshifts. Related interesting results were also obtained by Clark in her PhD [10] (see also [11]).
@@
Let $f\in\mathcal{L}$. Then exactly one of the following conditions hold:
@@ 2.1. Lorenz maps
As we see from the definition of Lorenz map $f$, it has discontinuities. It is also not completely clear if we should define $f(c)=0$ or $f(c)=1$, since both assignments are equally reasonable. To deal with this problem, we will use standard doubling points construction (see e. g. [35] for details) changing phase space $[0,1]$ into Cantor set $\mathbb{X}$, and $f$ into continuous map $\hat{f}$. In this construction all elements in $\left(\bigcup_{n=0}^{\infty}f^{-n}\{c\}\right)\setminus\{0,1\}$ are doubled (we perform a kind of Denjoy extension). We easily see that this new space differs from the original interval $[0,1]$ by at most countably many points and we do not modify endpoints, hence clearly the new space $\mathbb{X}$ is a Cantor set. If $I_{e}$ is inserted ,,hole” in place of a point $e$, and since $f$ preserves orientation of $[0,1]$, when defining $\hat{f}\colon\mathbb{X}\to\mathbb{X}$, we send its endpoints into respective endpoints of the hole $I_{f(e)}$, if there is a hole related to $f(e)$ or into the point $f(e)$ if it was not blown up (i.e. when it is 00 or $1$). Also, if we denote by $I_{c}=[c_{-},c_{+}]$ the hole related to critical point, and $f(0)=c$ then we define $\hat{f}(0)=c_{+}$, and the case $f(1)=c$ is dealt analogously. It is done the same way, when $f^{n}(0)=c$ but this time $\hat{f}(0)=a_{+}$ where $I_{f(0)}=[a_{-},a_{+}]$. Note that in this case also $a_{-}$ will be in image of $\hat{f}$, because $f$ is expanding and so in particular $f(1)>f(0)$. We have also a natural continuous projection $\pi\colon\mathbb{X}\to[0,1]$ “gluing” inserted points back together.
@@ 2.3. Lorenz maps and renormalization
(2.4) $\hat{F}_{g}=\{x\in\mathbb{X}:\operatorname{Orb}(x)\cap(\hat{f}^{k}(c_{+}),\hat{f}^{m}(c_{-}))\neq\emptyset\},$
@@
(2): Condition (2a) is a straightforward consequence of the definition. Condition (2b) is a simple consequence of monotonicity. To see this, fix any $x\in[0,u)$, $y\in(v,1]$ and denote $i=N((0,x))$ and $j=N((y,1))$. By the definition of $N(U)$ we obtain that $f(x)>c>u$ and $f(y)<c<v$. But if $x=0$ then similar argument to the above yields that $f^{k-1}(0)=u$ and the proof is completed by (1).
@@
Next, fix any $\hat{x}\in[0,\hat{u})$. Then $x=\pi(\hat{x})\in[u,1]$. But then (2b) implies that $\operatorname{Orb}(u)\cap(u,1]\neq\emptyset$. If $\operatorname{Orb}(\hat{u})\cap\{c_{-},c_{+}\}=\emptyset$ then we have $\operatorname{Orb}(\hat{u})\cap(\hat{u},1]\neq\emptyset$. But $\{c_{-},c_{+}\}\subset(\hat{u},1])$ so $\operatorname{Orb}(\hat{u})\cap(\hat{u},1]\neq\emptyset$ also when $\operatorname{Orb}(\hat{u})\cap\{c_{-},c_{+}\}\neq\emptyset$. Since $\hat{F}_{g}$ is completely invariant, we obtain $\hat{F}_{g}=\mathbb{X}$. The case $f(v)\in(u,v)$ is dealt the same way. The proof is completed. ∎
@@
Before proceeding further, let us recall Example 5.1 form [33].
@@
This conditions lead to the equations
@@
$\frac{1}{\beta+\beta^{2}}\approx 0.4299\leq\alpha\leq\frac{-\beta^{3}+\beta^{2}+2\beta-1}{\beta+\beta^{2}}\approx 0.4651.$
@@
$f^{4}((u,\hat{u}))=(\hat{u},\hat{v})\quad\text{ and }\quad f^{2}(\hat{u})\in(\hat{u},\hat{v}).$
@@
$P_{R}:=\min\{x\in O:x\geq c_{+}\}\quad\text{and}\quad\widetilde{P}_{R}:=\min\{x\in\widetilde{O}:x\geq c_{+}\}.$
@@
are well-defined. Moreover, by Theorem 4.7 we have $\hat{P}_{L}$, $\hat{P}_{R}\in\bigcup_{i=0}^{\infty}\hat{f}^{i}(U)$. First, we will prove that there are $l$, $r\in\mathbb{N}_{0}$ such that the set $\hat{f}^{l}(U)$ contains a left-sided neighbourhood of $\hat{P}_{L}$ and the set $\hat{f}^{r}(U)$ contains a right-sided neighbourhood of $\hat{P}_{L}$, i.e. $(a,\hat{P}_{L}]\subset\hat{f}^{l}(U)$ and $[\hat{P}_{L},b)\subset\hat{f}^{r}(U)$ for some points $a<\hat{P}_{L}<b$. Suppose it is not the case. In particular $\hat{P}_{L}\not\in\bigcup_{n=0}^{\infty}\operatorname{int}\hat{f}^{n}(U)$, hence $\hat{P}_{L}\in\operatorname{Orb}(\hat{f}(0))\cup\operatorname{Orb}(\hat{f}(1))$. Assume that there is a point $b>\hat{P}_{L}$ such that $[\hat{P}_{L},b)\subset\hat{f}^{r}(U)$ for some $r\in\mathbb{N}_{0}$, but for any $i\in\mathbb{N}_{0}$ and $a<\hat{P}_{L}$ there is a point $x\in(a,\hat{P}_{L})$ such that $x\notin\hat{f}^{i}(U)$. Note that in the case $\hat{P}_{L}\not\in\operatorname{Orb}(\hat{f}(0))$, if we pick $z_{0}\in U$ such that $\hat{f}^{i}(z_{0})=\hat{P}_{L}$ then for some $e$ sufficiently close to $z_{0}$ we have $(e,z_{0}]\subset U$ and $(\hat{f}^{i}(e),\hat{P}_{L}]=(\hat{f}^{i}(e),\hat{f}^{i}(z_{0})]\subset\hat{f}^{i}(U)$ which is a contradiction, therefore $\hat{P}_{L}\in\operatorname{Orb}(\hat{f}(0))$. We claim that $\hat{f}(0)$ is periodic. If we assume the contrary then there is a point $x_{0}\in(c_{+},1)$ such that $\hat{f}(x_{0})=\hat{f}(0)$ and $x_{0}\not\in\operatorname{Orb}(\hat{f}(0))$. Then $x_{0}\in\operatorname{Orb}(\hat{f}(1))$ because otherwise $x_{0}\in\operatorname{int}\hat{f}^{j}(U)$ for some $j$, and since we are not passing through any of $c_{+},c_{-}$, this gives $\hat{P}_{L}\in\operatorname{int}\hat{f}^{n}(U)$ for some $n$ which is a contradiction. But $x_{0}\in\operatorname{Orb}(\hat{f}(1))$ is a contradiction as well, since $\hat{f}(1)$ is periodic while $x_{0}$ cannot. Indeed, the claim holds and $\hat{f}(0)$ is periodic, which gives $\operatorname{Orb}(\hat{f}(0))=\hat{O}$. The symmetric case $(a,\hat{P}_{L}]\subset\hat{f}^{l}(U)$ leads to $\operatorname{Orb}(\hat{f}(1))=\hat{O}$.
@@
In the [7], the authors raise the question about parameters $\beta$ for which generalized $\beta$-transformation has matching (this class contain all linear Lorenz maps). They ask whether there exist generalised $\beta$-transformations with non-Pisot and non-Salem number $\beta$ and matching (see end of p.72 in [7]). Our Example 3.1 answers this question affirmatively, showing that such map exists, since map constructed there has matching by Corollary 4.15.
@@
It may happen that exactly one of the following conditions hold
@@
$Orb_{n}(x)\cap I_{i}\neq\emptyset\quad\text{for}\quad i=-1,0,1,\ldots,n-k-2,n-k,\ldots,n-1,$
@@ 6. Basic renormalizations
Let $f_{0}\colon[0,1]\to[0,1]$ be an expanding Lorenz map and $n\in\mathbb{N}$. Assume that there is a sequence of Lorenz maps $\{f_{i}\}_{i=1}^{n}$ such that for every $i=1,2,\ldots,n$ the map $f_{i}\colon[u_{i},v_{i}]\to[u_{i},v_{i}]$ has one of the following forms:
(6.1) $f_{i}=(f_{i-1},f_{i-1}^{2}),\quad f_{i}=(f_{i-1}^{2},f_{i-1})\quad\text{or}\quad f_{i}=(f_{i-1}^{2},f_{i-1}^{2}).$
Then for every $i=1,2,\ldots,n$ there exist numbers $l_{i}$, $r_{i}\in\mathbb{N}$ such that $f_{i}=(f_{0}^{l_{i}},f_{0}^{r_{i}})$. Moreover, the following trichotomy holds.

### v4

@@ 1. Introduction
The main object of our study is expanding Lorenz maps, that is maps $f\colon[0,1]\to[0,1]$ satisfying the following three conditions:
@@
with special emphasis on piecewise linear case. It turned out that such maps appear in a natural way as Poincarè maps in geometric models of well known Lorenz attractor. This description was independently appeared in works of Guckenheimer [21], Williams [37] and Afraimovich, Bykov and Shil‘nikov [4] as a tool for better understanding of chaotic behavior present in Lorenz model (cf. more recent [27]). In 1985, Hofbauer performed in [22] an analysis of possible periods of periodic points in Lorenz maps by introducing special type of infinite graphs, induced by evolution of pieces of monotonicity under iteration of Lorenz map (see [23] for further extensions of this approach). While a characterization similar to Scharkovsky was provided later in [3] by different methods, the tools developed by Hofbauer are of independent interest, since they provide a deep insight into dynamics of Lorenz maps. An extensive study on utility of Hofbauer’s approach, kneading theory of Milnor and Thurston and many other interesting results on dynamics of Lorenz maps can be found in PhD thesis of M. St. Pierre, see [32] (cf. [26]). Even the simplest case of maps $x\mapsto\beta x$ generates many difficulties and leads to deep results. It is usually performed at the level of symbolic dynamics, so-called $\beta$-shifts. In [31], Schmeling showed that many of these maps do not have the specification property, and later Thompson [36] provided nice representations of these systems together with their further descriptions (see [12] for recent advances that originated from this, together with a nice survey on studies on structure of entropy and topological pressure). Mixing in $x\mapsto\beta x+\alpha$ also reveals complicated structure of parameters [33] (cf. [14] for measure-theoretic mixing in a similar family of maps).
@@
    (1.1) $R_{E}f(x)=\begin{cases}f^{l}(x),&x\in[f^{r}(c_{+}),c)\\ f^{r}(x),&x\in(c,f^{l}(c_{-})]\\ \end{cases}$
@@
Theorem 1.1 is aiming at providing a beautiful connection between renormalizations and completely invariant sets. We are going to show that unfortunately, this result is not true as stated, however we are able to provide some possible alternatives (cf. Lemma 2.9 or Example 3.9). This way we show that some connections exist, however their complete description in full generality of Lorenz maps may be a challenging problem. At this point it is worth to mention closely related recent studies of R. Alcazar Barrera [1] (see also [2]), initiated in his PhD thesis, on invariant sets arising from omitting gaps is subshifts. Related interesting results were also obtained by Clark in her PhD [10] (see also [11]).
@@
Let $f\in\mathcal{L}$. Then exactly one of the following conditions holds:
@@ 2.1. Lorenz maps
As we see from the definition of Lorenz map $f$, it has discontinuities. It is also not completely clear if we should define $f(c)=0$ or $f(c)=1$, since both assignments are equally reasonable. To deal with this problem, we will use standard doubling points construction (see e. g. [35] for details) changing phase space $[0,1]$ into Cantor set $\mathbb{X}$, and $f$ into continuous map $\hat{f}$. In this construction all elements in $\left(\bigcup_{n=0}^{\infty}f^{-n}\{c\}\right)\setminus\{0,1\}$ are doubled (we perform a kind of Denjoy extension). We easily see that this new space differs from the original interval $[0,1]$ by at most countably many points and we do not modify endpoints, hence clearly the new space $\mathbb{X}$ is a Cantor set. If $I_{e}$ is inserted ”hole” in place of a point $e$, and since $f$ preserves orientation of $[0,1]$, when defining $\hat{f}\colon\mathbb{X}\to\mathbb{X}$, we send its endpoints into respective endpoints of the hole $I_{f(e)}$, if there is a hole related to $f(e)$ or into the point $f(e)$ if it was not blown up (i.e. when it is 00 or $1$). Also, if we denote by $I_{c}=[c_{-},c_{+}]$ the hole related to critical point, and $f(0)=c$ then we define $\hat{f}(0)=c_{+}$, and the case $f(1)=c$ is dealt analogously. It is done the same way, when $f^{n}(0)=c$ but this time $\hat{f}(0)=a_{+}$ where $I_{f(0)}=[a_{-},a_{+}]$. Note that in this case also $a_{-}$ will be in image of $\hat{f}$, because $f$ is expanding and so in particular $f(1)>f(0)$. We have also a natural continuous projection $\pi\colon\mathbb{X}\to[0,1]$ “gluing” inserted points back together.
@@ 2.3. Lorenz maps and renormalization
(2.4) $\hat{F}_{g}=\{x\in\mathbb{X}:\operatorname{Orb}(x)\cap(\hat{f}^{k}(c_{+}),\hat{f}^{m}(c_{-}))\neq\emptyset\}.$
@@
(2): Condition (2a) is a straightforward consequence of the definition. Condition (2b) is a simple consequence of monotonicity. To see this, fix any $x\in(0,u)$, $y\in(v,1)$ and denote $i=N((0,x))$ and $j=N((y,1))$. By the definition of $N(U)$ we obtain that $f^{i}(x)>c>u$ and $f^{j}(y)<c<v$. But if $x=0$ then similar argument to the above yields that $f^{k-1}(0)=u$ and the proof is completed by (1).
@@
Next, fix any $\hat{x}\in[0,\hat{u})$. Then $x=\pi(\hat{x})\in[0,u]$. But then (2b) implies that $\operatorname{Orb}(u)\cap(u,1]\neq\emptyset$. If $\operatorname{Orb}(\hat{u})\cap\{c_{-},c_{+}\}=\emptyset$ then we have $\operatorname{Orb}(\hat{u})\cap(\hat{u},1]\neq\emptyset$. But $\{c_{-},c_{+}\}\subset(\hat{u},1]$ so $\operatorname{Orb}(\hat{u})\cap(\hat{u},1]\neq\emptyset$ also when $\operatorname{Orb}(\hat{u})\cap\{c_{-},c_{+}\}\neq\emptyset$. Since $\hat{F}_{g}$ is completely invariant, we obtain $\hat{F}_{g}=\mathbb{X}$. The case $f(v)\in(u,v)$ is dealt the same way. The proof is completed. ∎
@@
Before proceeding further, let us recall Example 5.1 from [33].
@@
These conditions lead to the equations
@@
$\frac{1}{\beta+\beta^{2}}\approx 0.4299\leq\alpha\leq\frac{-\beta^{3}+\beta^{2}+2\beta-1}{\beta+\beta^{2}}\approx 0.4651,$
@@
$f^{4}((u,\hat{u}))=(\hat{u},\hat{v})\quad\text{ and }\quad f^{2}(\hat{u})\in(\hat{u},\hat{v}),$
@@
$P_{R}:=\min\{x\in O:x\geq c_{+}\}\quad\text{and}\quad\widetilde{P}_{R}:=\min\{x\in\widetilde{O}:x\geq c_{+}\},$
@@
are well-defined. Moreover, by Theorem 4.7 we have $\hat{P}_{L}$, $\hat{P}_{R}\in\bigcup_{i=0}^{\infty}\hat{f}^{i}(U)$. First, we will prove that there are $l$, $r\in\mathbb{N}_{0}$ such that the set $\hat{f}^{l}(U)$ contains a left-sided neighborhood of $\hat{P}_{L}$ and the set $\hat{f}^{r}(U)$ contains a right-sided neighborhood of $\hat{P}_{L}$, i.e. $(a,\hat{P}_{L}]\subset\hat{f}^{l}(U)$ and $[\hat{P}_{L},b)\subset\hat{f}^{r}(U)$ for some points $a<\hat{P}_{L}<b$. Suppose it is not the case. In particular $\hat{P}_{L}\not\in\bigcup_{n=0}^{\infty}\operatorname{int}\hat{f}^{n}(U)$, hence $\hat{P}_{L}\in\operatorname{Orb}(\hat{f}(0))\cup\operatorname{Orb}(\hat{f}(1))$. Assume that there is a point $b>\hat{P}_{L}$ such that $[\hat{P}_{L},b)\subset\hat{f}^{r}(U)$ for some $r\in\mathbb{N}_{0}$, but for any $i\in\mathbb{N}_{0}$ and $a<\hat{P}_{L}$ there is a point $x\in(a,\hat{P}_{L})$ such that $x\notin\hat{f}^{i}(U)$. Note that in the case $\hat{P}_{L}\not\in\operatorname{Orb}(\hat{f}(0))$, if we pick $z_{0}\in U$ such that $\hat{f}^{i}(z_{0})=\hat{P}_{L}$ then for some $e$ sufficiently close to $z_{0}$ we have $(e,z_{0}]\subset U$ and $(\hat{f}^{i}(e),\hat{P}_{L}]=(\hat{f}^{i}(e),\hat{f}^{i}(z_{0})]\subset\hat{f}^{i}(U)$ which is a contradiction, therefore $\hat{P}_{L}\in\operatorname{Orb}(\hat{f}(0))$. We claim that $\hat{f}(0)$ is periodic. If we assume the contrary then there is a point $x_{0}\in(c_{+},1)$ such that $\hat{f}(x_{0})=\hat{f}(0)$ and $x_{0}\not\in\operatorname{Orb}(\hat{f}(0))$. Then $x_{0}\in\operatorname{Orb}(\hat{f}(1))$ because otherwise $x_{0}\in\operatorname{int}\hat{f}^{j}(U)$ for some $j$, and since we are not passing through any of $c_{+},c_{-}$, this gives $\hat{P}_{L}\in\operatorname{int}\hat{f}^{n}(U)$ for some $n$ which is a contradiction. But $x_{0}\in\operatorname{Orb}(\hat{f}(1))$ is a contradiction as well, since $\hat{f}(1)$ is periodic while $x_{0}$ cannot. Indeed, the claim holds and $\hat{f}(0)$ is periodic, which gives $\operatorname{Orb}(\hat{f}(0))=\hat{O}$. The symmetric case $(a,\hat{P}_{L}]\subset\hat{f}^{l}(U)$ leads to $\operatorname{Orb}(\hat{f}(1))=\hat{O}$.
@@
In the [7], the authors raise the question about parameters $\beta$ for which generalized $\beta$-transformation has matching (this class contain all linear Lorenz maps). They ask whether there exist generalized $\beta$-transformations with non-Pisot and non-Salem number $\beta$ and matching (see end of p.72 in [7]). Our Example 3.1 answers this question affirmatively, showing that such map exists, since map constructed there has matching by Corollary 4.15.
@@
It may happen that exactly one of the following conditions holds
@@
$Orb_{n}(x)\cap I_{i}\neq\emptyset\quad\text{for}\quad i=-1,0,1,\ldots,n-k-2,n-k,\ldots,n-1.$
@@ 6. Basic renormalizations
Let $f\colon[0,1]\to[0,1]$ be an expanding Lorenz map and $n\in\mathbb{N}$. Assume that there is a sequence of Lorenz maps $\{f_{i}\}_{i=1}^{n}$ such that for every $i=1,2,\ldots,n$ the map $f_{i}\colon[u_{i},v_{i}]\to[u_{i},v_{i}]$ has one of the following forms:
(6.1) $f_{i}=(f_{i-1},f_{i-1}^{2}),\quad f_{i}=(f_{i-1}^{2},f_{i-1})\quad\text{or}\quad f_{i}=(f_{i-1}^{2},f_{i-1}^{2}),$
where $f_{0}:=f$. Then for every $i=1,2,\ldots,n$ there exist numbers $l_{i}$, $r_{i}\in\mathbb{N}$ such that $f_{i}=(f_{0}^{l_{i}},f_{0}^{r_{i}})$. Moreover, the following trichotomy holds.
@@ Acknowledgements
Publication co-financed by the Krakow University of Economics as part of the “Conference Activity Support program – WAK-2024”.